# Beyond Silent Letters: Amplifying LLMs in Emotion Recognition with Vocal Nuances

### v1

@@ Abstract
This paper introduces a novel approach to emotion detection in speech using Large Language Models (LLMs). We address the limitation of LLMs in processing audio inputs by translating speech characteristics into natural language descriptions. Our method integrates these descriptions into text prompts, enabling LLMs to perform multimodal emotion analysis without architectural modifications. We evaluate our approach on two datasets: IEMOCAP and MELD, demonstrating significant improvements in emotion recognition accuracy, particularly for high-quality audio data. Our experiments show that incorporating speech descriptions yields a 2 percentage point increase in weighted F1 score on IEMOCAP (from 70.111% to 72.596%). We also compare various LLM architectures and explore the effectiveness of different feature representations. Our findings highlight the potential of this approach in enhancing emotion detection capabilities of LLMs and underscore the importance of audio quality in speech-based emotion recognition tasks. We’ll release the source code on Github.
@@ 2.1 LLM for emotion recognition
Similar extensions to InstructERC have been explored in recent literature. BiosERC Xue et al. (2024) enhances the model by extracting and injecting "biographical information" of speakers into the system, demonstrating improved performance across multiple datasets. CKERC Fu (2024) takes a different approach, incorporating commonsense knowledge about interlocutors’ reactions and intentions to deepen the understanding of speaker information. Both these methods, like our approach, aim to enrich the input to LLMs with additional contextual information.
@@ 2.2 Speech Incorporation in LLM
The integration of speech features into LLM-based systems for emotion recognition and related tasks has been explored in various recent works, each offering unique approaches and insights. SECap Xu et al. (2023) introduces a framework that uses LLaMA as a text decoder to generate coherent speech emotion captions instead of labels. While SECap demonstrates the potential of LLMs in describing speech emotions, it differs from our approach by requiring an additional audio encoder (HuBERT) and a bridging network (Q-Former) to process speech inputs. Similarly, Zhang et al. (2024) proposes an approach to integrate speech into LLMs for depression detection by adding acoustic landmark tokens to the input prompt. This method shares our goal of incorporating speech information directly into LLM inputs but necessitates additional parameters and pre-training steps to introduce audio information.
@@ 4.2 Primary Results
Our results from Table 1 demonstrate that incorporating speech descriptions into the text input yields performance improvements across both datasets, albeit to varying degrees. For the IEMOCAP dataset, we observe a substantial improvement of nearly 2 percentage points when adding speech descriptions to the text input (from 70.111% to 72.021%). By providing additional speech features for context, the model improves to 72.596%. (Detailed experiments are in Appendix A.1) This significant gain underscores the value of integrating acoustic information in emotion recognition tasks, particularly for high-quality audio recordings.
In contrast, the MELD dataset shows a more modest improvement with the addition of speech impression, increasing from 67.44% to 67.604%. This disparity in performance gain between the two datasets can be attributed to the differences in audio quality and recording conditions. IEMOCAP, recorded in a controlled studio environment, provides clean and consistent audio, allowing for more accurate extraction of speech features. On the other hand, MELD, derived from the TV show "Friends," presents more challenging audio conditions. The presence of background laughter, varying audio quality, overlapping speakers, and short utterance durations (averaging 2-3 seconds) in MELD likely contribute to less reliable speech feature extraction, thereby limiting the potential benefits of speech descriptions.

### v2

@@ Abstract
This paper introduces a novel approach to emotion detection in speech using Large Language Models (LLMs). We address the limitation of LLMs in processing audio inputs by translating speech characteristics into natural language descriptions. Our method integrates these descriptions into text prompts, enabling LLMs to perform multimodal emotion analysis without architectural modifications or addition. We evaluate our approach on two datasets: IEMOCAP and MELD, demonstrating significant improvements in emotion recognition accuracy, particularly for high-quality audio data. Our LLM-finetuning experiments show that incorporating speech descriptions yields a more than 2 percentage point increase in average weighted F1 score on IEMOCAP (from 70.111% to 72.596%). We also compare various LLM architectures and explore the effectiveness of different feature representations. Our findings highlight the potential of this approach in enhancing emotion detection capabilities of LLMs and underscore the importance of audio quality in speech-based emotion recognition tasks. We’ll release the source code on Github.
@@ 2.1 LLM for emotion recognition
Similar extensions to InstructERC have been explored in recent literature. BiosERC Xue et al. (2024) enhances the model by extracting and injecting "biographical information" of speakers into the system, demonstrating improved performance across multiple datasets. CKERC Fu (2024) takes a different approach, incorporating commonsense knowledge about interlocutors’ reactions and intentions to deepen the understanding of speaker information. DialogueLLM Zhang et al. (2024b) incorporates visual descriptions generated from GPT-4 into the prompt. All of these methods, like our approach, aim to enrich the input to LLMs with additional contextual information.
@@ 2.2 Speech Incorporation in LLM
The integration of speech features into LLM-based systems for emotion recognition and related tasks has been explored in various recent works, each offering unique approaches and insights. SECap Xu et al. (2023) introduces a framework that uses LLaMA as a text decoder to generate coherent speech emotion captions instead of labels. While SECap demonstrates the potential of LLMs in describing speech emotions, it differs from our approach by requiring an additional audio encoder (HuBERT) and a bridging network (Q-Former) to process speech inputs. Similarly, Zhang et al. (2024a) proposes an approach to integrate speech into LLMs for depression detection by adding acoustic landmark tokens to the input prompt. This method shares our goal of incorporating speech information directly into LLM inputs but necessitates additional parameters and pre-training steps to introduce audio information.
@@ 4.2 Primary Results
Our results from Table 1 demonstrate that incorporating speech features into the text input yields performance improvements across both datasets, albeit to varying degrees. For the IEMOCAP dataset, we observe a substantial improvement of nearly 2 percentage points when adding speech descriptions to the text input (from 70.111% to 72.021%). By providing additional speech features for context, the model improves to 72.596%. (Detailed experiments are in Appendix A.1) This significant gain underscores the value of integrating acoustic information in emotion recognition tasks, particularly for high-quality audio recordings.
In contrast, the MELD dataset does not show improvemnet using speech descriptions and a modest improvement with the addition of speech impression, increasing from 67.44% to 67.604%. This disparity in performance between the two datasets can be attributed to the differences in audio quality and recording conditions. IEMOCAP, recorded in a controlled studio environment, provides clean and consistent audio, allowing for more accurate extraction of speech features. On the other hand, MELD, derived from the TV show "Friends," presents more challenging audio conditions. The presence of background laughter, varying audio quality, overlapping speakers, and short utterance durations (averaging 2-3 seconds) in MELD likely contribute to less reliable speech feature extraction, thereby limiting the potential benefits of speech descriptions.
@@ A.1 Speech features for context
In our exploration of incorporating speech features into context utterances, we found that selective application of these features yielded better results than a blanket approach. As shown in Table 4, adding speech descriptions to only the last three utterances in the context proved more effective than applying them to all utterances. This suggests that recent context carries more weight in emotion recognition. Notably, pitch features outperformed volume features, with pitch information from the last three utterances providing a 0.6% improvement over non-contextual speech descriptions, achieving a weighted F1 score of 72.595%. Interestingly, combining all speech features for context utterances did not enhance performance, likely due to information overload. These findings highlight the importance of judicious feature selection and context consideration in multimodal emotion recognition tasks.